How do you write an equation for a line that is parallel to the given line and passes through the given point? This is for the equation: y = 5x-8; (3,5)
Answer by ankor@dixie-net.com(22740)

How do you write an equation for a line that is parallel to the given line and passes through the given point?
This is for the equation: y = 5x-8; (3,5)
:
We use the point/slope form: y - y1 = m(x - x1)
Parallel lines have the same slope, therefore:
m = 5; x1=3, y1=5
y - 5 = 5(x - 3)
y - 5 = 5x - 15
y = 5x - 15 + 5
y = 5x - 10, is the equation